OpenStudy (anonymous):

each year a school organizes a concert. in 2004 the cost of organizing the concert was $385. in 2005 the cost was 10% less than in 2004. calculate the cost in 2005.

OpenStudy (lgbasallote):

\[$385 - ($385 \times 0.10)\]

OpenStudy (lgbasallote):

need i say more?

OpenStudy (anonymous):

wait theres another part for this question

OpenStudy (anonymous):

why 0.10?

OpenStudy (lgbasallote):

10% = 0.10

OpenStudy (anonymous):

i know but why did we convert it? and are u sure of this answer?

OpenStudy (lgbasallote):

very

OpenStudy (lgbasallote):

it says that in 2005 it is 10% less...so we must look for what is 10% of the money in 2004

OpenStudy (lgbasallote):

the money in 2004 is 385 10% of 385 = 385 x 0.10

OpenStudy (lgbasallote):

now it says less so we subtract that from the money in 2004 so it becomes 385 - (385 x 0.10)

OpenStudy (anonymous):

ok ima tell you part 2

OpenStudy (lgbasallote):

shoot

OpenStudy (anonymous):

the cost of $385 in 2004 was 10% more than the cost in 2003. calculate the cost in 2003

OpenStudy (lgbasallote):

wait i analyzed it wrong...

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

$385−($385×0.10) =$346.5 detail sol the cost in 2004=$385 a/c to ques cost in 2005 is less than 10 % of385=$38.5 cost in 2005=$385-$38.5=$346.5

OpenStudy (lgbasallote):

wahhhh analyzed it wrong again =(( darn trick question :p haha

OpenStudy (anonymous):

:P

OpenStudy (lgbasallote):

oops lol...i was right =))) got confused and doubted myself lol 385 = a + (a* 0.10) 385 = a + 0.10a 385 = 1.10a can you solve for a on your own?

OpenStudy (anonymous):

naa, im too dumb for solvin math

OpenStudy (anonymous):

btw there are 3 more parts for it :P

OpenStudy (lgbasallote):

o.O

OpenStudy (anonymous):

yeah :P

OpenStudy (lgbasallote):

okay fine lay it on me

OpenStudy (anonymous):

:D

OpenStudy (lgbasallote):

btw the cost for 2003 is a = 385/1.10 let the calculator do the work there

OpenStudy (anonymous):

in 2006 the number of tickets sold was 210. the ratio; (no. of adults: no. of student tickets was 23:19. how many tickets were sold?

OpenStudy (anonymous):

no wait the ratio'swrong

OpenStudy (anonymous):

its no.of adult tickets:no.of student tickets

OpenStudy (lgbasallote):

this is a different question o.O

OpenStudy (anonymous):

no its bart b)

OpenStudy (anonymous):

ok :D

OpenStudy (lgbasallote):

yeah..close lol..

OpenStudy (lgbasallote):

btw...i notice you are new to the site? it is tradition to the site that we medal people who helpus to show our appreciation and to show that we got the concept ;D